# A flexible and general semi-supervised approach to multiple hypothesis testing

### v1

@@ 6.2 Real data experiments with p-values
We next evaluate RESET and related methods on a collection of publically available datasets with p-values that were investigated by Zhang et al. [50] when introducing AdaFDR and by Lei and Fithian [33] when introducing AdaPT. Since AdaFDR and AdaPT operate under essentially the same assumptions as RESET’s, we expect these datasets to therefore satisfy Assumption 3 with the exception of two datasets, which we discuss at the end of this section. We provide a brief summary of each dataset below:
@@
  • A proteomics dataset (Proteomics [14]): the goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells. Each protein consists of side information equal to the natural log of the number of peptides “quantified” across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
@@
  • fMRI datasets (Auditory and Imagination) [47]: the goal is to identify voxels that are activated in response to different stimuli. In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis. Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label [5], which is used to delineate different areas of the brain with different functions.
@@
  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): the goal is to identify single nucleotide polymorphisms (SNPs) that are associated with gene expression levels. Such an association is referred to as an expressive quantitative trait loci (eQTL). Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene trascription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
@@
  • Gene-drug response data (Estrogen [12]): the goal is to identify genes that are associated with a change in gene expression levels in response to a low-dosage estrogen treatment applied to breast cancer cells. Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment. Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2). Thus, we can analyze the data twice, once for each side information variable based on (1) and (2).
@@
Figure 12: Comparison of RESET and other methods using a collection of publically available datasets. Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. Each color and shape combination corresponds to a unique method. For readability, each point is jittered in the horizontal direction.
@@
The results of each method applied to the collection of the above datasets is visually summarized in Figure 12. We can clearly see that RESET Ensemble is comparable to the other methods in terms of discoveries. Indeed, to argue this more accurately, we ranked each method across each dataset and at each FDR threshold considered (ties were averaged). Then we calculated the median rank for each method. RESET Ensemble and AdaPTg-GMM were tied with the highest median ranks.
@@
There are a couple of implementation details we would like to highlight here (for full details, see Supplementary Sections S.5-S.10). First, we used the same default options of RESET Ensemble on all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1. In this case, it is sensible to define asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively as Chao and Fithian [8] did in their application of AdaPTg-GMM to this dataset. We also point out that we implemented AdaPTg-GMM and ZAP using the p-value information, and not the test statistic information (see Section 3.1 for background on these methods). To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics. For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d uniformly. In other words, we input plausible z-values that correspond to the same p-value information used by the other tools. As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
@@
To evaulate the computation times for each method, we used the two fMRI datasets, which uses the most number of side information variables and has a large number of hypotheses. Figure 13(A-B) shows the computation times for each considered method where the fastest are the collection of RESET methods, ZAP and AdaFDR. RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
@@
Noticeably, we have excluded the analysis of the two eQTL studies from Figure 12. This is because this dataset exhibits a high degree of p-value dependency, even among the true null hypotheses and thus violates Assumption 3 which guarantees our type-1 error control. Indeed, this phenomenon is due to linkage disequilibrium in which SNPs behave dependently. As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large suggesting a low confident hypothesis. Figure 13(C-D) shows that RESET Ensemble reports substantially more discoveries on these eQTL datasets than all the non-RESET variants. This is because the underlying random forest in the ensemble is able to identify the same ‘copies’ of low confident p-values in both the training and test sets, and if these copies are mutually positively labelled, which many are, then they will be rescored higher by our RESET algorithm and ultimately discovered. Evidently, RESET NN is more robust to this problem: the number of discoveries is in the same ‘ballpark’ as other methods. Note that RESET GAM shares identical results with RESET RF since it fails to perform a spline basis expansion on one of the side information variables (the chromatin state of the SNP), and so RESET GAM switches to RESET RF in this case.
@@ 6.3 FDP control with p-values
Figure 14: Comparison of RESET and other methods with FDP control using a collection of publically available data. Each panel plots the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
In this final section, we demonstrate RESET’s application of FDP control in the p-value setting. Here we define the following asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively, for all datasets. The rationale stems from our later analysis in the discussion, as well as our comments at the end of Section 5.2.3, which point out the fact that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power. Hence, to rectify this, we essentially double the number of estimating decoys under consideration by defining a decoy region that is twice as large as the corresponding target region.
@@
Figure 14 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section, with varying FDR thresholds (1%, 5%, 10%, 20%) and confidence levels (50%, 80%, 90%). RESET Ensemble is much more powerful than the alternative non-side information approaches with few exceptions.
@@ 7 Discussion
We introduced RESET, a flexible wrapper for any semi-supervised learning algorithm. Alongside, we provide an implementation of RESET, called RESET Ensemble, that (1) is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, (2) is fast compared to most tools, and (3) is able to achieve finite FDR/FDP control.
@@
In evaluating the claim of (1), we highlight the fact that competing tools such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants. Making a selection after observing the discovery lists of each tool will clearly cheat any guarantee of FDR control. Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or by the variants separately, which was the case of AdaPT. Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power to many of these tools. RESET Ensemble’s intended generality alleviates the user from having to choose the “right” method to begin with.
@@
In the case of (2), we found that our collection of RESET methods were a faster alternative to all competing methods with finite FDR control in our considered experiments. The only case in which RESET was surpassed in terms of speed was ZAP, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 13B). The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which the variants of Adaptive Knockoffs were generally too slow to be used as tools for this application.
@@
Finally in addressing (3), we point out that RESET’s compatibility with FDR or FDP control is unique. In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure. Katsevich et al. [27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds. However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDP at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDP control). To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b). Similairly, we know of no method that allows finite sample FDP control in the p-value setting with side information either.
@@
On the topic of RESET’s FDP control, we point out the following limitation. When the number of discoveries is relatively large, RESET yields more power over the generic FDP-SD. However, when the number of discoveries is small, or the confidence is too high, RESET may yield less power, even if the side information is reasonably informative. To illustrate this, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, that are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$). We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$. Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
@@
Figure 15: RESET’s and FDP-SD’s bounds on the number of decoy wins. We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$. On the right y-axis, we plot the ratio of these bounds (in black). The left panel looks at indices between 1K to 2K, while the right panel looks at indices from 10K to 20K. An index corresponds to the number of top scoring hypotheeses (regardless of their labels).
@@
Figure 15 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices. We can see that for smaller indices, this ratio is much smaller, initially at 50% at an index of 1K, while at 10K this ratio is much higher at 94%. If the order of the hypotheses is the same, RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t. $D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labelled hypotheses. In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K. If the side information is not rich enough, it is possible RESET will report few discoveries than FDP-SD in these cases. We intend on investigating this as part of future work.
@@
Figure 16: Variability in the number of discoveries in the Airway data and in Simulation 5 In (A) We used a 10% FDR threshold and varied the internal seed of RESET Ensemble over 100 applications. In (B) we used a 10% FDR threshold and compared the variability in RESET’s and AdaPT’s power. There are two outliers in RESET’s case (which occurred when the signal was considerably low).
@@
Notably, RESET’s discovery list is variable since it randomly splits the decoys into two sets: one set for training and the other for estimating the false discoveries. This variability is demonstrated in the histogram of the number of discoveries in Figure 16A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%. The resulting number of discoveries has a mean of 5998 with a standard deviation of about 85. While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times. Indeed, in Figure 16(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT’s even though AdaPT does not employ any internal randomization. In practice, users may get around RESET’s randomization by setting the RNG as a function of the input data and parameters. We plan on adding this feature to a future update.
@@
Lastly, while RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting. There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$. Moreover there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35]. Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications.
Appendix A Supplement
@@
4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy wins in top $k$ scores
5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/feate wins in top $k$ scores
@@
5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo targets
@@
where the first line follows since the denominator sums to 1, and the second line clearly follows from the non-decreasing property. For a more formal argument for the second line, note that the non-decreasing property implies that:
@@
where $B\in[0,a)$, $p_{i}$ is a true null p-value and $q$ denotes the inverse transformation of the mirrored p-values that map p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$. Then consider for every measurable-$A$:
@@
where $B$ is the corresponding region such that $\{W_{i}\in A\}=\{p_{i}\in B\}\cup\{p_{i}\in q(B)\}$. It follows that:
@@
$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i})$
@@
See 3
In this section we prove that given Assumption 3 RESET controls the FDR or FDP. We use the target-decoy terminology to refer to a positive label $L_{i}=1$ as a target win or simply a target, and a negative label $L_{i}=-1$ as a decoy win or decoy.
@@
Let $\tilde{L}_{i}$ for $i\in[p]$ denote the training labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo target. Let $\mathcal{F}$ be the $sigma$-algebra generated by the winning scores $W$, the side information $\mathbf{x}$, and the labels of the false nulls. Let $N$ be the indices of the true nulls, and $L_{-i},\tilde{L}_{-i}$ be the corresponding labels without the $i$th one. Then we have for $i\in N$:
$L_{i}$ =$=\begin{cases}+1,&\mathbb{P}(L_{i}=+1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})}\\ -1,&\mathbb{P}(L_{i}=-1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})}\\ \end{cases},$
independently of $\{\tilde{L}_{i}=1\},\mathcal{F},\tilde{L}_{-i},L_{-i}$.
@@
By Assumption 3, $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$. In addition, since each decoy is randomly assigned to the training decoy set independently of anything else with probability $s$, for each $i\in N$ we have:
$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\mathbb{P}(L_{i}=-1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1,\mathcal{F},\tilde{L}_{-i},L_{-i})$
@@
and
$\mathbb{P}(L_{i}=1,\tilde{L}_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})$ $=\mathbb{P}(L_{i}=1\mid\mathcal{F},\tilde{L}_{-i},L_{-i})\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=1,\mathcal{F},\tilde{L}_{-i},L_{-i})$
$\leq c_{0}\cdot 1,$

### v2

@@ 6.2 Real data experiments with p-values
We next evaluate RESET and related methods on a collection of publicly available datasets with p-values that were investigated by Zhang et al. [51] when introducing AdaFDR and by Lei and Fithian [33] when introducing AdaPT. Since AdaFDR and AdaPT operate under essentially the same assumptions as RESET’s, we expect these datasets to therefore satisfy Assumption 3 with the exception of two datasets, which we discuss at the end of this section. We provide a brief summary of each dataset below.
@@
  • A proteomics dataset (Proteomics [14]): The goal is to identify proteins that have different protein abundances when treated with rapamycin versus DMSO in yeast cells. Each protein consists of side information equal to the natural log of the number of peptides ‘quantified’ across all samples, i.e., the number of peptides belonging to the protein with their abundances measured.
@@
  • Two fMRI datasets (Auditory and Imagination) [48]: The goal is to identify voxels that are activated in response to different stimuli. In the first dataset, a single individual received auditory stimulus and in the second dataset, a single individual was instructed to imagine playing tennis. Each hypothesis (voxel) is equipped with four dimensional side information: the spatial position of the voxel (in three dimensions) and a categorical variable with the Brodmann label [5], which is used to delineate different areas of the brain with different functions.
@@
  • Two eQTL datasets (Subcutaneous and Omentum [11, 10]): The goal is to identify single nucleotide polymorphisms (SNPs) that are associated with changes in gene expression levels. Such an association is referred to as an expression quantitative trait locus (eQTL). Each hypothesis (SNP) is equipped with the following four side information variables: (1) the distance between the SNP and the gene transcription’s start site (TSS), (2) the log gene expression level, (3) the alternative allele frequency (AAF) of the SNP and (4) the chromatin state of the SNP.
@@
  • Gene-drug response data (Estrogen [12, 36]): The goal is to identify genes that are differentially expressed in response to a low-dosage estrogen treatment applied to breast cancer cells. Each hypothesis (gene) is assigned a ‘rank’ that was generated using an analysis of the gene expression levels in response to (1) a high-dosage estrogen treatment and (2) a medium-dosage estrogen treatment. Here, the ranks are in terms of the strength of association between the expression level of each gene and the level of dosage in (1) and (2). Thus, we can analyze the data twice, once for each side information variable based on (1) and (2).
@@
Figure 12: Comparison of RESET and other methods using a collection of publicly available datasets. Each method was evaluated at the FDR thresholds of 1%, 5%, 10%, and 20%. The number of discoveries reported by each RESET method is averaged over 10 applications at FDR threshold and dataset. Each color and shape combination corresponds to a unique method. For readability, each point is jittered in the horizontal direction.
@@
The results of each method applied to the collection of the above datasets is visually summarized in Figure 12. We can clearly see that RESET Ensemble is comparable to the other methods in terms of discoveries. Indeed, to argue this more accurately, we ranked each method across each dataset and at each FDR threshold considered (ties were averaged). Then we calculated the median rank for each method. RESET Ensemble and AdaPTg-GMM were tied with the highest median rank. Moreover, for each dataset and each FDR threshold in Figure 12, we computed the maximum number of discoveries across all methods and determined the relative power of each method to this maximum. The boxplots in Figure 13 display the relative powers for each method. Visually, RESET Ensemble and AdaPTg-GMM have the highest medians, with RESET Ensemble at 95.9% and AdaPTg-GMM at 94.5% with RESET Ensemble’s boxplot edging higher overall.
Figure 13: Relative power of each method. Each boxplot describes the relative power of each method computed for each dataset and FDR threshold in Figure 12.
@@
There are a couple of implementation details we would like to highlight here (for full details, see Supplementary Sections S.6-S.10). First, we used the same default options of RESET Ensemble on all datasets, except for the two fMRI datasets which exhibit a spike of p-values close to 1. In this case, it is sensible to define asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively as Chao and Fithian [8] did in their application of AdaPTg-GMM to this dataset. We also point out that we applied AdaPTg-GMM and ZAP using the p-value information rather than the test statistic information (see Section 3.1 for background on these methods). To do so, AdaPTg-GMM allows the user to directly input p-values instead of z-statistics. For ZAP, we converted the p-value information into z-values by defining $z_{i}=\pm\Phi^{-1}(p_{i}/2)$ of each p-value, $p_{i}$, where the signs are chosen i.i.d. uniformly. In other words, we input plausible z-values that correspond to the same p-value information used by the other tools. As we mentioned in Section 3, our implementation of RESET focuses on the use of p-values, and we leave the extension to test statistic information as part of future work.
@@
To evaluate the computation times for each method, we used the two fMRI datasets, which uses the largest number of side information variables and has a large number of hypotheses. Figure 14(A-B) shows the computation times for each considered method, where the fastest are the collection of RESET methods, ZAP and AdaFDR. RESET NN, RESET GAM and RESET RF are slightly faster than our ensemble approach which considers all of them.
@@
Notably, we have excluded the analysis of the two eQTL studies from Figure 12. This is because this dataset exhibits a high degree of p-value dependency, even among the true null hypotheses and thus violates Assumption 3 which guarantees our type-1 error control. Indeed, this phenomenon is due to linkage disequilibrium in which SNPs behave dependently. As a consequence, the dataset exhibits virtually identical copies of the same p-value, with similar side information, even if the p-value is large, suggesting a true null hypothesis. Figure 14(C-D) shows that RESET Ensemble reports substantially more discoveries on these eQTL datasets than all the non-RESET variants. This is because many estimating and training decoys are virtually the same and the random forest is sensitive to this leakage of information. Evidently, RESET NN is more robust to this problem: the number of discoveries is in the same ‘ballpark’ as other methods. Note that RESET GAM shares identical results with RESET RF since it fails to perform a spline basis expansion on one of the side information variables (the chromatin state of the SNP), and so RESET GAM switches to RESET RF in this case.
@@ 6.3 FDP control with p-values
Figure 15: Comparison of RESET and other methods with FDP control using a collection of publicly available data. Each panel plots the number of discoveries obtained by RESET Ensemble, FDP-SD and GR-SD at a range of FDR thresholds (1%, 5%, 10%, 20%) with varying confidence $1-\gamma$ (0.5, 0.8, 0.9).
In this final section, we demonstrate RESET’s application of FDP control in the p-value setting. Here we define the following asymmetric target and decoy regions, $[0,0.3)$ and $(0.3,0.9]$ respectively, for all datasets. The rationale stems from our later analysis in the Discussion, as well as our comments at the end of Section 5.2.3, which point out that when the ratio of estimating decoys to true null targets is small, the subsequent FDP-SD step in RESET might incur a loss of power. Hence, to rectify this, we essentially double the number of estimating decoys under consideration by defining a decoy region that is twice as large as the corresponding target region.
@@
Figure 15 displays the number of discoveries obtained by each method on the collection of real datasets we considered in the previous section, with varying FDR thresholds (1%, 5%, 10%, 20%) and confidence levels (50%, 80%, 90%). RESET Ensemble is much more powerful than the alternative non-side information approaches with a few exceptions.
@@ 7 Discussion
We introduced RESET, a flexible wrapper for any semi-supervised learning algorithm. Alongside, we provide an implementation of RESET, called RESET Ensemble, that is power-wise competitive with recently developed methods for competition as well as p-value based multiple testing problems with side information, fast compared to most tools, and flexible as it allows finite-sample of either FDR or FDP control.
@@
In terms of power, we highlight the fact that competing tools, such as the variants of AdaPT or Adaptive Knockoffs, require an initial selection of one of those variants. Making a selection after observing the discovery lists of each tool will generally compromise its FDR control. Thus, we compared RESET Ensemble to the default competing method if one was available, which was the case of Adaptive Knockoff’s GAM filter, or separately to each variant, which was the case of AdaPT. Based on our comparisons, we conclude that RESET Ensemble is able to achieve comparable or slightly greater power than many of these tools while freeing the user from having to choose the ‘right’ method to begin with.
@@
As far as computational runtime, we found that our collection of RESET methods provide a faster alternative to all competing methods with finite FDR control in our considered experiments. The only case in which RESET was surpassed in terms of speed was ZAP-asymp, which demonstrated marginally faster runtimes than RESET Ensemble in one of the fMRI datasets (Figure 14B). However, ZAP-asymp’s fast runtimes come at the cost of asymptotic FDR control. The most striking difference in runtimes was in the peptide detection context using mass spectrometry data, in which all variants of Adaptive Knockoffs were generally too slow to be practical.
@@
Finally, we point out that RESET’s compatibility with FDR or FDP control is unique. In the competition setting, we note that it would be challenging to combine Adaptive Knockoffs with FDP-SD because FDP-SD is a stepdown procedure while Adaptive Knockoffs is essentially a step-up procedure. Katsevich and Ramdas [27] provide a general method for adaptive ordering using their method for simultaneous FDP bounds. However, (a) they do not provide an exact implementation of this, only a theoretical discussion, and (b) their approach is conservative since it controls the FDP at every cutoff (which allows the analyst to explore the nested discovery lists without breaking FDP control). To the best of our knowledge, in the competition framework with side information, RESET is the only method of reordering that may be flexibly used with FDP-SD without the additional cost of (b). Similarly, we know of no method that allows finite sample FDP control in the p-value setting with side information.
@@
On the topic of RESET’s FDP control, we point out the following limitation. When the number of discoveries is small, e.g., when the confidence is too high, RESET may yield less power, even if the side information is reasonably informative. To illustrate this phenomenon, we computed the bounds, $\delta_{i}$ on the number of decoy wins in the top $i$ scores from Algorithm S2, which are used to determine the reported list of discoveries in FDP-SD (with $c=1/2$) and RESET (using the default $c=2/3$). We considered a confidence level of $1-\gamma=90\%$ and an FDP threshold of $\alpha=1\%$. Since we removed approximately half the decoys for training purposes in RESET, the bounds need to be doubled to fairly compare between the two approaches.
@@
Figure 16: RESET’s and FDP-SD’s bounds on the number of decoy wins. We recorded, on the left y-axis, the bounds used to determine the list of FDP-SD discoveries and twice the bounds used to determine RESET’s discoveries at a range of indices at $\alpha=1\%$ and $1-\gamma=90\%$. On the right y-axis, we plot the ratio of these bounds (in black). The left panel looks at indices between 501 to 2K, while the right panel looks at indices from 10K to 20K. An index corresponds to the number of top scoring hypotheses (regardless of their labels).
@@
Figure 16 shows the bounds for FDP-SD, and double the bounds for RESET, along with the ratio of the two at a range of different indices. We can see that for smaller indices, this ratio is much smaller, initially at 0% at an index of 501 and 50% at an index of 1K, while at 10K this ratio is much higher at 94%. If the order of the hypotheses is the same, then RESET’s smaller bounds imply it will report fewer discoveries because RESET employs FDP-SD to control the FDP by searching for the largest index $i$ s.t. $D_{i_{0}}\leq\delta_{i},\dots,D_{i}\leq\delta_{i}$, and reports the top $i$ scoring positively labeled hypotheses. In other words, if RESET is to report roughly the same number of discoveries as FDP-SD, then RESET needs to rearrange the targets and estimating decoys successfully enough so as to make up for a 50% difference in the bounds when considering the top 1K scores, while only 6% when considering the top 10K. If the side information is not rich enough, then it is possible RESET will report few discoveries than FDP-SD in these cases. We intend to investigate this as part of future work.
@@
Figure 17: Variability in the number of discoveries. in the Airway data and in Simulation 5 (A) Histogram of the number of RESET Ensemble discoveries applied 100 times to the Airway data with a 10% FDR threshold and varying its internal seed. (B) Boxplots of RESET’s and AdaPT’s power in 100 applications of the ‘circle in the middle’ setup of Simulation 5 using a 10% FDR threshold. There are two outliers in RESET’s case (which occurred when the signal was considerably low).
@@
Notably, RESET’s discovery list is variable since it randomly splits the decoys into two sets: one set for training and the other for estimating the false discoveries. This variability is demonstrated in the histogram of the number of discoveries in Figure 17A using 100 applications of RESET Ensemble to the Airway dataset at an FDR threshold of 10%. The resulting number of discoveries has a mean of 5998 with a standard deviation of about 85. While this is problematic to an extent, keep in mind that the data distribution is random, and the added variance from RESET may be marginal at times. Indeed, in Figure 17(B), we find that in the case of Simulation 5 from Section 6.1.1, RESET is overall arguably less variable than AdaPT, even though AdaPT does not employ any internal randomization. In practice, users may get around RESET’s randomization by setting the random number generator to be a function of the input data and parameters. We plan on adding this feature to a future update.
@@
Lastly, while RESET is a flexible method for the competition and p-value based testing, we wish to further extend it to the test-statistic setting. There is more than one way of generalizing RESET to accommodate a test statistic $z_{i}$ for each hypothesis $H_{i}$. Moreover, there are several ‘types’ of interactions between the side information $\mathbf{x}_{i}$ and the test statistic $z_{i}$ which are not preserved when switching to p-values, as outlined by Leung and Sun [35]. Thus it is not clear that RESET Ensemble will be equally applicable in this setting without substantial modifications.
Supplement
@@
4:$D_{k}\leftarrow\#\{i\leq k:L_{i}=-1\}$triangleright number of decoy/knockoffs wins in top $k$ scores
5:$T_{k}\leftarrow\#\{i\leq k:L_{i}=1\}$triangleright number of target/original feature wins in top $k$ scores
@@
5:$\tilde{L}_{i}\leftarrow 1$ for $i\in J:=I^{c}$triangleright pseudo-targets
@@
where the first line follows since the denominator sums to 1, and the second line clearly follows from the non-decreasing property as explained next.
First, note that the non-decreasing density property implies that:
@@
where $B\subseteq[0,a)$ is measurable, $p_{i}$ is a true null p-value and $q$ denotes the inverse transformation of the mirrored p-values that map p-values in $(b_{1},b_{2}]$ onto $[0,a)$, that is $q(p):=b_{2}-\frac{b_{2}-b_{1}}{a}\cdot p$.
Let $A\subseteq\mathbb{R}$ be measurable and let $B\subseteq[0,a)$ be such that $\{W_{i}\in A\}=\{p_{i}\in B\}\cup\{p_{i}\in q(B)\}$. Then,
@@
$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in q(B))$
[…]
It follows that:
@@
$\leq\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W_{i})=\frac{a}{b_{2}-b_{1}}\mathbb{P}(p_{i}\in(b_{1},b_{2}]\mid W,\mathbf{x},L_{-i}).$
@@
In this section we prove that given Assumption 3 RESET controls the FDR or the FDP. We use the target-decoy terminology to refer to a positive label $L_{i}=1$ as a target win or simply a target, and a negative label $L_{i}=-1$ as a decoy win or decoy.
@@
Let $\tilde{L}_{i}$ for $i\in[p]$ denote the pseudo labels, where $\tilde{L}_{i}=-1$ for a training decoy and $\tilde{L}_{i}=1$ for a pseudo-target. Let $\mathcal{F}$ be the $sigma$-algebra generated by the winning scores $W$, the side information $\mathbf{x}$, and the labels of the false nulls. Let $N$ be the indices of the true nulls, and $L_{-i},\tilde{L}_{-i}$ be the corresponding labels without the $i$th one. Then under Assumption 3, we have for $i\in N$, $L_{i}=\pm 1$ with
$\mathbb{P}(L_{i}=+1\mid\tilde{L}_{i}=1)\leq\frac{c_{0}}{1-s\cdot(1-c_{0})},\quad\mathbb{P}(L_{i}=-1\mid\tilde{L}_{i}=1)\geq\frac{(1-c_{0})\cdot(1-s)}{1-s\cdot(1-c_{0})},$
independently of $\mathcal{F},\tilde{L}_{-i},L_{-i}$.
@@
By Assumption 3, we know $\{L_{i}\,:\,i\in N\}$ are i.i.d with $P(L_{i}=1)\leq c_{0}$, independently of $\mathcal{G}:=\sigma(\mathcal{F},\tilde{L}_{-i},L_{-i})$. Hence, the following probability statements are independent of $\mathcal{G}$. In addition, since each decoy is randomly assigned to the training decoy set independently of anything else with probability $s$, for each $i\in N$ we have:
$\mathbb{P}(L_{i}=-1,\tilde{L}_{i}=1)$ $=\mathbb{P}(L_{i}=-1)\cdot\mathbb{P}(\tilde{L}_{i}=1\mid L_{i}=-1)$